Teal Mountain Inc. had a beginning inventory of 98 units of Product RST at a cost of $9 per unit. During the year, purchases were:

Feb. 20

557 units at $10

Aug. 12

410 units at $12

May 5

470 units at $11

Dec. 8

95 units at $13

Teal Mountain Inc. uses a periodic inventory system. Sales totaled 1,350 units.
Determine the ending inventory and the cost of goods sold under each of the assumed cost flow methods (FIFO, LIFO, and average-cost). (Round answers to 0 decimal places, e.g. 150.)

FIFO

LIFO

AVERAGE-COST

The ending inventory

$enter a dollar amount

Entry field with incorrect answer now contains modified data

$enter a dollar amount

Entry field with incorrect answer

$enter a dollar amount

Entry field with incorrect answer

The cost of goods sold

$enter a dollar amount

Entry field with incorrect answer

$enter a dollar amount

Entry field with incorrect answer

$enter a dollar amount

Entry field with incorrect answer
Answer #1

X fac units beginning inventory 20-Feb 05-May 12-Aug 08-Dec units 98 557 470 410 95 1,630 1,350 280 cost per unit $9 $ 10 $ 1

for formulas and calculations, refer to the image below -

se units units units cost per unit total cost =C3*D3 =C4*D4 =C5*D5 =C6*D6 =C7*D7 =SUM(E3:E7) beginning inventory 98 43516 557

In case you have any query, kindly ask in comments.